# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ 5.2 Literature Comparisons
In Tables 5.4, 5.4, and 5.4 we list the currently known protostructures in the three high C3VO fields as reported in the literature. These lists contain only protostructures that have been spectroscopically confirmed to some level, either through galaxy-traced methods or through intergalactic medium (IGM) tomographic methods. In addition to the spectroscopically-confirmed protostructures reported in the literature, we also include all spectroscopic overdensity candidates from the VUDS survey that were found via the spectroscopic overdensity methods described in Lemaux et al. (2014a) and Lemaux et al. (2018). These candidates are also referred to as protostructures for the remainder of the paper for simplicity. Central coordinates and redshifts are taken directly from the literature for literature-reported structures and from a unit-weighted positional mean of all potential member galaxies for VUDS spectroscopic candidate overdensities. For protostructures reported in C3VO studies (Cucciati et al., 2014, 2018; Lemaux et al., 2018; Shen et al., 2021; Forrest et al., 2023; Shah et al., 2024a; Staab et al., 2024), we list the galaxy overdensity-weighted barycenter associated with each peak of a given protostructure when available. The various peaks are denoted either by the number appended to each structure name or by the proper name of the peak. The numbers of each peak match those in the referenced study. The references associated with each protostructure are listed in the tables.
@@
As we did for the matching process between simulated structure and candidates in the mock observations of the lightcone, we search for candidates for each protostructure listed in Tables 5.4, 5.4, and 5.4 within a line of sight distance of and a transverse radius of d h Mpc. Since many of these protostructures are observed over much larger transverse scales, and since some protostructures (e.g., those from Newman et al. 2022) have fairly large positional uncertainties, we additionally search on scales of d h Mpc if no associated candidates are found within d h Mpc (though still employing a line of sight criterion of ). If multiple candidates are found within the search volume, the candidate with the largest associated mass is reported. We note that, beyond initial testing of the overdensity search algorithm using the VUDS candidate overdensities and a few of the published C3VO protostructures, the search algorithm was not optimized to find overdensities in the literature. Further, with the exception of a few pointed C3VO observations in the ECDFS field, in which we included DEIMOS/MOSFIRE observations targeting the protostructure “Smruti” at , and in the COSMOS field, in which we included two DEIMOS masks targeting the protostructure “Taralay” at , we do not include any spectroscopic observations specifically targeting any of the protostructures included in this comparison in the construction of our maps. The inclusion of such observations would only serve to increase our recovery fraction.
@@
In total, 97 protostructures are listed in Tables 5.4, 5.4, and 5.4. Of these, 74% (72/97) have a match to a candidate in our catalog within and d h Mpc and 91% (86/97) have a match within and d h Mpc. Of the 14 protostructures that have no associated candidate(s) within d h Mpc but do have associated candidates within the larger radial search of d h Mpc, a large fraction (40%) are of peaks in known multicomponent systems (e.g., Hyperion, Elentári, PClJ0227-0421) where blending of the signal from the various peaks is likely to be an issue for our search algorithm. These blending issues are also likely to be at least partially responsible for the non-detections associated with four of the peaks of the known C3VO protostructures. We emphasize here that the search algorithm developed in this paper is optimized to detect overdensities at high redshift, not to parse out their morphology or distribution of substructure at a detailed level. Once structure is detected, methods such as those used in Cucciati et al. (2018), Shen et al. (2021), Forrest et al. (2023), Shah et al. (2024a), and Staab et al. (2024) can be used to further zoom in on a given protostructure and map its density distribution on a finer scale. Of the nine protostructures that have no associated candidate(s), four (LATIS-D2-1, QPC.28, PG.4.53, and Tosh20.2) lie at the boundaries of our maps in the transverse or line of sight dimensions.
@@
As a reminder, most of the redshifts obtained from Keck/DEIMOS and Keck/MOSFIRE as part of the dedicated followup of protostructures in the three C3VO fields were not used in the creation of our maps. The exceptions are those redshifts described in Shen et al. (2021), Lemaux et al. (2022), and Shah et al. (2024a), which comprise only 25% of the full C3VO-Keck spectroscopic database and are almost exclusively limited to three protostructures (Smruti in ECDFS, Taralay in COSMOS, and PClJ027-0421 in CFHTLS-D1). Additionally, no redshifts obtained as part of the dedicated followup of any of the literature protostructures discussed in this section were used in the making of our density maps.
@@
Overall, our mapping and search algorithms show an impressive recovery of known protostructures. We note that the protostructures listed only constitute roughly 20% of the candidates presented in our catalog. Figure 14 shows a comparison between the redshifts and Gaussian amplitudes for previously known protostructures versus those that are newly discovered in this work. As expected, the previously known structures are primarily concentrated among the candidates with the largest Gaussian amplitudes (which are presumably the most massive candidates in these fields) as can be seen in the approximately exponentially increasing fraction of known structures with detection strength across all fields. We are able to find lower strength candidates in all three fields up to about , with the redshift boundary in CFHTLS-D1 appearing especially sharp. Due to our lack of completeness at higher redshifts, we are unable to conclusively comment on the behavior seen in these plots for candidates at .
@@ 5.3 Comparison to the LATIS Survey
Of note is the efficacy of the search algorithm in recovering signal for massive protostructures in the Ly Tomography IMACS Survey (LATIS; Newman et al., 2020, 2022) via intergalactic medium (IGM) tomography. This survey covers three fields, two of which overlap with C3VO: the CFHTLS-D1 field and the COSMOS field (i.e., CFHTLS-D2). The eight protostructures reported in Newman et al. (2022), of which we compare to seven in our study101010The eighth protostructure is in the CFHTLS-D4 field, a field which is not covered by C3VO., are expected from simulations to be very massive () but are generally not well traced by rest-frame UV-selected galaxies. Despite an apparent paucity of such galaxies in many of these protostructures, six out of the seven strongest absorbing systems in LATIS as reported in Newman et al. (2022) are recovered in our search. The only undetected protostructure (LATIS-D2-1) is near the edge of our VMC map and is likely prevented from being detected due to border masking. However, despite having successfully detected protostructures associated with most of the strongest LATIS absorption systems, the strength of the C3VO galaxy-traced candidates are moderate in general, with a median Gaussian amplitude of 200 for the detected LATIS structures versus a median Gaussian amplitude of 420 for all recovered structures taken from the literature.
@@
The results presented in Newman et al. (2022) were based on a partial spectroscopic data set with only a portion of the planned survey depth, breadth, and sampling rate. The survey has since been completed and IGM maps have been recomputed in all three fields covered by the survey (A. Newman, priv. communication). For brevity, we refer to the second generation maps, those computed with the full survey data, as “LATIS2” and the first generation maps that formed the basis of Newman et al. (2022) as “LATIS1”. The strongest large-scale IGM detections in the LATIS2 maps are reported in Newman et al. (in prep). This sample includes the 19 strongest detections in COSMOS and the eight strongest detections in the CFHTLS-D1 field. Of these detections, we find counterparts for 17 of the 19 (89%) in the COSMOS field and four of the eight (50%) in the CFHTLS-D1 field within Mpc and 11111111 of the 20 counterparts are found within Mpc and .. Only one of the detections in the LATIS2 maps, LATIS2-D2-01 (formerly LATIS1-D2-1), lies at the edge of our maps and is likely undetectable. All other detections are well contained within the areal and redshift coverage of our maps. Interestingly, we do not detect any associated structure in our maps for either of the two strongest detections in the LATIS2 map in the CFHTLS-D1 field. This absence, and the generally decreased efficacy of our method at detecting LATIS structures in the CFHTLS-D1 field is likely due to the decreased quality of data in that field in terms of the spectroscopic sampling, the photometric redshift uncertainties, and the shallower NIR data. The average Gaussian amplitude for LATIS2 counterparts detected in our maps is 210, slightly higher than that of the LATIS1 counterparts (200), but only roughly half the strength of that of our typical literature counterpart (420).
@@
Since our density mapping is based on a combined spectroscopic and photometric sample that is limited in the rest-frame optical, making it effectively a stellar mass limited sample, the decrease in strength of LATIS detections may indicate that such structures are not only deficient in rest-frame UV selected galaxies but rather in galaxies in general. However, the relative contribution of the spectroscopic sample in our mapping decreases with increasingly red colors, where redder star-forming galaxies and quiescent galaxies primarily enter the mapping through their photometric redshifts, which complicates interpretation. Ultimately, a full census in the rest-frame optical is necessary to make any definitive conclusions about the overall paucity of galaxies in LATIS structures.
@@ 5.4 Comparison to the ODIN Survey
A comparison between C3VO overdensities at in the COSMOS field, using maps generated in a similar manner to what was used here, and ODIN precursor NB422 narrowband observations taken with the One-Degree Imager mounted on the Wisconsin-Indiana-Yale-NOIRLab (WIYN) Observatory targeting LAEs at the same redshift was already presented in Huang et al. (2022). These observations broadly mimic those of ODIN, though they are slightly shallower, target only the lowest redshift that ODIN targets (), and are taken in a filter with a FWHM roughly twice as wide as that of the ODIN observations (75 Å vs. 170Å). In this section we concentrate on a comparison between protostructure candidate detections in C3VO and those detected in ODIN in the COSMOS field at . Specifically, we focus on maps generated by Ramakrishnan et al. (2023) of the extended COSMOS field from early release data. These maps cover nearly an order of magnitude larger area than the 1.3 deg2 C3VO maps in the COSMOS field. A comparison of overdensities between the two surveys at and in the COSMOS field, as well as the comparisons in other fields will be the subject of future work.
@@
As discussed extensively in Huang et al. (2022), LAEs appear to trace overdensities in a complicated manner. While certain structures appear to be well traced by LAEs, with some structures exhibiting higher levels of LAEs overdensity relative to other galaxy types, other structures appear to be largely devoid of LAEs (e.g., Apostolovski et al., 2024). In the Hyperion proto-supercluster (Cucciati et al., 2018), Huang et al. (2022) found that LAEs generally trace the overall galaxy density distribution well, with several of the most prominent matter peaks in Hyperion exhibiting strong LAE overdensities, and a significant correlation was observed between the density distribution of LAEs and spectroscopically-confirmed member galaxies. However, at least one of the peaks (peak 5, Wang et al. 2016) was not well traced by LAEs, perhaps due, at least in part, to advanced evolutionary stage of that peak. Similar variation between the density of LAEs and the overall galaxy population has also been observed in several case studies of overdensities at high redshift (e.g., Shi et al., 2019a, b) and potentially originates from variations in the density of large-scale, cool Hi gas reservoirs, which have the effect of diffusing or, in the presence of dust, destroying Ly photons, causing an overall suppression of LAEs. Indeed, Huang et al. (2022) also found, by cross-correlating IGM tomographic maps from the COSMOS Lyman-Alpha Mapping And Mapping Observations (CLAMATO, Lee et al. 2016, 2018) and LAEs, that while LAEs appear to trace structure moderately rich in large-scale, cool Hi gas well, LAEs tend to avoid regions with the highest Hi column densities (see also Momose et al. 2021). A multitude of other considerations could also modulate the overdensity of LAEs within large-scale structures, including the prevalence of outflows due to potential increased AGN activity (e.g., Shah et al., 2024b) or star-formation activity (e.g., Lemaux et al., 2022; Staab et al., 2024), changes in the age and/or dust content of constituent member galaxies, and small number statistics due in part to LAEs being fairly rare populations at these redshifts (e.g., Cassata et al., 2015).
@@
In Figure 15 we show a comparison between ODIN-detected protoclusters and C3VO protostructure candidates in the region of COSMOS where the ODIN maps and our VMC maps overlap. We limit the C3VO protostructure candidates to those with Gaussian amplitudes 100 and a redshift range . The amplitude cut was made in order to include only those C3VO candidates with a 50% chance of being a genuine overdensity. The redshift range was chosen to order to roughly match redshift selection function of ODIN N501 narrowband observations () with an additional () buffer in order to account for uncertainties associated with C3VO systemic redshifts and the wings of the NB501 throughput curve. In total, three ODIN protoclusters lie within the overlap volume, all three of which have at least one C3VO protostructure candidate whose barycenter falls within a projected distance of 2.5 Mpc of the ODIN protocluster barycenters. Both ODIN protoclusters in the northeast of the overlap volume appear to be potentially associated with multiple C3VO protostructure candidates at a variety of different redshifts. The ODIN protocluster at [RA, Dec] = [150.54, 2.52] also appears to be associated with MAGAZ3NEJ1001, which is an overdensity that, perhaps not surprisingly given its presence in the ODIN map, appears to consist of primarily bluer star-forming galaxies, demographics that hold even for its most massive members (McConachie et al., 2024). All C3VO protostructure candidates that are within a proximity of 2.5 Mpc of ODIN protoclusters have moderate Gaussian amplitudes in the range of .
@@
In contrast, the inverse comparison reveals that the two most strongly detected, and presumably most massive, protostructure candidates in the C3VO map in this volume (), which are located at [RA, Dec] = [150.25, 2.25], do not have counterpart overdensities in the ODIN maps. Additionally, several protostructure candidates with Gaussian amplitudes that are similar to those associated with the ODIN protoclusters do not have associated detections in the ODIN maps, with % of C3VO candidates lying further than 2.5 Mpc from an ODIN protocluster. However, three out of the four of the C3VO candidates that are not associated with an ODIN protocluster lie either outside or at the edge of the redshift selection window for ODIN LAEs. The one C3VO candidate with a systemic redshift that lies well within the ODIN redshift selection window (located at [RA, Dec] = 150.36, 2.35) has a barycenter that lies within 5 Mpc of an ODIN-selected protocluster in projection. Furthermore, the two strongest C3VO detections in this volume were detected as potential overdensities in Ramakrishnan et al. (2023), but did not exhibit a strong enough overdensity over a sufficiently large area to pass their fiducial cuts for classification as a genuine protocluster. As both of these C3VO candidates lie at the extrema of the ODIN redshift selection window, it is likely that the ODIN observations are only able to recover a small fraction of member galaxies in these systems.
@@
Structure | RAlit | Declit | RAC3VO | DecC3VO | Amp.aaAmplitude of the Gaussian fit to the linked candidate | dtrans | IDC3VO | ref. | ||
@@
C.PS4 | 150.4500 | 2.3710 | 2.035 | 150.4395 | 2.3726 | 2.024 | 265 | 0.33 | 1 | 1,2 |
COSTCO2 | 149.8710 | 2.2290 | 2.047 | 149.9323 | 2.1311 | 2.089 | 162 | 3.57∗ | 228 | 3 |
ZFIRE | 150.0940 | 2.2510 | 2.095 | 150.0459 | 2.2415 | 2.112 | 322 | 1.51+ | 336 | 4 |
G237 | 150.5070 | 2.3120 | 2.160 | 150.4889 | 2.3617 | 2.146 | 237 | 1.62 | 429 | 5,6 |
COSTCO3 | 150.1290 | 2.2750 | 2.160 | 150.0957 | 2.2449 | 2.208 | 284 | 1.38 | 666 | 3 |
C.PS6 | 150.1900 | 2.1330 | 2.174 | 150.1786 | 2.1887 | 2.153 | 166 | 1.74 | 433 | 1,2 |
C.PS8 | 149.8640 | 2.0900 | 2.216 | 149.8603 | 2.0950 | 2.215 | 615 | 0.19 | 770 | 1,2 |
BD2.2 | 150.1970 | 2.0030 | 2.232 | 150.1848 | 2.0176 | 2.226 | 929 | 0.58+ | 715, 791 | 7 |
C.PS9 | 149.9200 | 2.2210 | 2.277 | 149.8851 | 2.1652 | 2.284 | 455 | 2.00 | 1129 | 1,2,8 |
COSTCO5 | 149.9380 | 2.0910 | 2.283 | 1 50.0112 | 2.1181 | 2.262 | 98 | 2.38 | 985 | 3 |
COSTCO1 | 150.1100 | 2.1610 | 2.298 | 150.1021 | 2.2218 | 2.301 | 380 | 1.86 | 1054, 1156 | 3 |
COSTCO4 | 149.7060 | 2.0240 | 2.391 | 149.7715 | 2.1389 | 2.416 | 305 | 3.99∗,+ | 1626 | 3 |
Theia | 150.0937 | 2.4049 | 2.468 | 150.0951 | 2.4086 | 2.468 | 988 | 0.12+ | 1918, 2142 | 9, 10, 11 |
Eos | 149.9765 | 2.1124 | 2.426 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.31 | 1685 | 9 |
Helios | 149.9996 | 2.2537 | 2.444 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.25 | 1685 | 8, 9, 12, 13 |
Selene | 150.2556 | 2.3423 | 2.469 | 150.1992 | 2.3781 | 2.462 | 2013 | 2.00+ | 1866 | 9, 11 |
Hyp5 | 150.2293 | 2.3381 | 2.507 | 150.1992 | 2.3781 | 2.462 | 2013 | 1.50+ | 1886 | 9, 14 |
Hyp6 | 150.3316 | 2.2427 | 2.492 | 150.3403 | 2.2236 | 2.496 | 370 | 0.63+ | 1982 | 9 |
Hyp7 | 149.9581 | 2.2187 | 2.423 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.01 | 1685 | 9 |
LATIS1-D2-0 | 150.1030 | 2.1750 | 2.562 | 150.1327 | 2.3380 | 2.558 | 417 | 4.93∗,+ | 2291 | 15 |

### v2

@@ 5.2 Literature Comparisons
In Tables 5.4, 5.4, and 5.4 we list the currently known protostructures in the three high- C3VO fields as reported in the literature. These lists contain only protostructures that have been spectroscopically confirmed to some level, either through galaxy-traced methods or through intergalactic medium (IGM) tomographic methods. In addition to the spectroscopically confirmed protostructures reported in the literature, we also include all spectroscopic overdensity candidates from the VUDS survey that were found via the spectroscopic overdensity methods described in Lemaux et al. (2014a) and Lemaux et al. (2018). These candidates are also referred to as protostructures for the remainder of the paper for simplicity. Central coordinates and redshifts are taken directly from the literature for literature-reported structures and from a unit-weighted positional mean of all potential member galaxies for VUDS spectroscopic candidate overdensities. For protostructures reported in C3VO studies (Cucciati et al., 2014, 2018; Lemaux et al., 2018; Shen et al., 2021; Forrest et al., 2023; Shah et al., 2024a; Staab et al., 2024), we list the galaxy overdensity-weighted barycenter associated with each peak of a given protostructure when available. The various peaks are denoted either by the number appended to each structure name or by the proper name of the peak. The numbers of each peak match those in the referenced study. The references associated with each protostructure are listed in the Tables.
@@
As we did for the matching process between simulated structure and candidates in the mock observations of the lightcone, we search for candidates for each protostructure listed in Tables 5.4, 5.4, and 5.4 within a line-of-sight distance of and a transverse radius of d h Mpc. Since many of these protostructures are observed over much larger transverse scales, and since some protostructures (e.g., those from Newman et al. 2022) have fairly large positional uncertainties, we additionally search on scales of d h Mpc if no associated candidates are found within d h Mpc (though still employing a line-of-sight criterion of ). If multiple candidates are found within the search volume, the candidate with the largest associated mass is reported. We note that, beyond initial testing of the overdensity search algorithm using the VUDS candidate overdensities and a few of the published C3VO protostructures, the search algorithm was not optimized to find overdensities in the literature. Further, with the exception of a few pointed C3VO observations in the ECDFS field, in which we included DEIMOS/MOSFIRE observations targeting the protostructure “Smruti” at , and in the COSMOS field, in which we included two DEIMOS masks targeting the protostructure “Taralay” at , we do not include any spectroscopic observations specifically targeting any of the protostructures included in this comparison in the construction of our maps. The inclusion of such observations would only serve to increase our recovery fraction.
@@
In total, 97 protostructures are listed in Tables 5.4, 5.4, and 5.4. Of these, 74% (72/97) have a match to a candidate in our catalog within and d h Mpc and 91% (86/97) have a match within and d h Mpc. Of the 14 protostructures that have no associated candidate(s) within d h Mpc but do have associated candidates within the larger radial search of d h Mpc, a large fraction (40%) are of peaks in known multicomponent systems (e.g., Hyperion, Elentári, PClJ0227-0421) where blending of the signal from the various peaks is likely to be an issue for our search algorithm. These blending issues are also likely to be at least partially responsible for the nondetections associated with four of the peaks of the known C3VO protostructures. We emphasize here that the search algorithm developed in this paper is optimized to detect overdensities at high redshift, not to parse out their morphology or distribution of substructure at a detailed level. Once structure is detected, methods such as those used in Cucciati et al. (2018), Shen et al. (2021), Forrest et al. (2023), Shah et al. (2024a), and Staab et al. (2024) can be used to further zoom in on a given protostructure and map its density distribution on a finer scale. Of the nine protostructures that have no associated candidate(s), four (LATIS-D2-1, QPC.28, PG.4.53, and Tosh20.2) lie at the boundaries of our maps in the transverse or line-of-sight dimensions.
@@
As a reminder, most of the redshifts obtained from Keck/DEIMOS and Keck/MOSFIRE as part of the dedicated follow-up of protostructures in the three C3VO fields were not used in the creation of our maps. The exceptions are those redshifts described in Shen et al. (2021), Lemaux et al. (2022), and Shah et al. (2024a), which comprise only 25% of the full C3VO-Keck spectroscopic database and are almost exclusively limited to three protostructures (Smruti in ECDFS, Taralay in COSMOS, and PClJ027-0421 in CFHTLS-D1). Additionally, no redshifts obtained as part of the dedicated follow-up of any of the literature protostructures discussed in this Section were used in the making of our density maps.
@@
Overall, our mapping and search algorithms show an impressive recovery of known protostructures. We note that the protostructures listed only constitute roughly 20% of the candidates presented in our catalog. Figure 14 shows a comparison between the redshifts and Gaussian amplitudes for previously known protostructures versus those that are newly discovered in this work. As expected, the previously known structures are primarily concentrated among the candidates with the largest Gaussian amplitudes (which are presumably the most massive candidates in these fields), as can be seen in the approximately exponentially increasing fraction of known structures with detection strength across all fields. We are able to find lower strength candidates in all three fields up to about , with the redshift boundary in CFHTLS-D1 appearing especially sharp. Due to our lack of completeness at higher redshifts, we are unable to conclusively comment on the behavior seen in these plots for candidates at .
@@ 5.3 Comparison to the LATIS Survey
Of note is the efficacy of the search algorithm in recovering signal for massive protostructures in the Ly Tomography IMACS Survey (LATIS; Newman et al., 2020, 2022) via IGM tomography. This survey covers three fields, two of which overlap with C3VO: the CFHTLS-D1 field and the COSMOS field (i.e., CFHTLS-D2). The eight protostructures reported in Newman et al. (2022), of which we compare to seven in our study999The eighth protostructure is in the CFHTLS-D4 field—a field that is not covered by C3VO., are expected from simulations to be very massive () but are generally not well traced by rest-frame UV-selected galaxies. Despite an apparent paucity of such galaxies in many of these protostructures, six out of the seven strongest absorbing systems in LATIS as reported in Newman et al. (2022) are recovered in our search. The only undetected protostructure (LATIS-D2-1) is near the edge of our VMC map and is likely prevented from being detected due to its proximity to the region where we consider our density reconstruction to be reliable. However, despite having successfully detected protostructures associated with most of the strongest LATIS absorption systems, the strength of the C3VO galaxy-traced candidates is moderate in general, with a median Gaussian amplitude of 200 for the detected LATIS structures versus a median Gaussian amplitude of 420 for all recovered structures taken from the literature.
@@
The results presented in Newman et al. (2022) were based on a partial spectroscopic data set with only a portion of the planned survey depth, breadth, and sampling rate. The survey has since been completed, and IGM maps have been recomputed in all three fields covered by the survey (A. Newman 2025, private communication). For brevity, we refer to the second-generation maps, those computed with the full survey data, as “LATIS2” and the first-generation maps that formed the basis of Newman et al. (2022) as “LATIS1”. The strongest large-scale IGM detections in the LATIS2 maps are reported in Newman et al. (2025, in preparation). This sample includes the 19 strongest detections in COSMOS and the eight strongest detections in the CFHTLS-D1 field. Of these detections, we find counterparts for 17 of the 19 (89%) in the COSMOS field and four of the eight (50%) in the CFHTLS-D1 field within Mpc and 101010Eleven of the 20 counterparts are found within Mpc and .. Only one of the detections in the LATIS2 maps, LATIS2-D2-01 (formerly LATIS1-D2-1), lies at the edge of our maps and is likely undetectable. All other detections are well contained within the areal and redshift coverage of our maps. Interestingly, we do not detect any associated structure in our maps for either of the two strongest detections in the LATIS2 map in the CFHTLS-D1 field. This absence, and the generally decreased efficacy of our method at detecting LATIS structures in the CFHTLS-D1 field, is likely due to the decreased quality of data in that field in terms of the spectroscopic sampling, the photometric redshift uncertainties, and the shallower NIR data. The average Gaussian amplitude for LATIS2 counterparts detected in our maps is 210, slightly higher than that of the LATIS1 counterparts (200), but only roughly half the strength of that of our typical literature counterpart (420).
@@
Since our density mapping is based on a combined spectroscopic and photometric sample that is limited in the rest-frame optical, making it effectively a stellar-mass-limited sample, the decrease in strength of LATIS detections may indicate that such structures are not only deficient in rest-frame UV-selected galaxies but rather in galaxies in general. However, the relative contribution of the spectroscopic sample in our mapping decreases with increasingly red colors, where redder star-forming galaxies and quiescent galaxies primarily enter the mapping through their photometric redshifts, which complicates interpretation. Ultimately, a full census in the rest-frame optical is necessary to make any definitive conclusions about the overall paucity of galaxies in LATIS structures.
@@ 5.4 Comparison to the ODIN Survey
A comparison between C3VO overdensities at in the COSMOS field, using maps generated in a similar manner to what was used here, and ODIN precursor NB422 narrowband observations taken with the One-Degree Imager mounted on the Wisconsin–Indiana–Yale–NOIRLab (WIYN) Observatory targeting LAEs at the same redshift that was already presented in Huang et al. (2022). These observations broadly mimic those of ODIN, though they are slightly shallower, target only the lowest redshift that ODIN targets (), and are taken in a filter with an FWHM roughly twice as wide as that of the ODIN observations (75 Å versus 170Å). In this Section, we concentrate on a comparison between protostructure candidate detections in C3VO and those detected in ODIN in the COSMOS field at . Specifically, we focus on maps generated by Ramakrishnan et al. (2023) of the extended COSMOS field from early release data. These maps cover nearly an order-of-magnitude larger area than the 1.3 deg2 C3VO maps in the COSMOS field. A comparison of overdensities between the two surveys at and in the COSMOS field, as well as the comparisons in other fields will be the subject of future work.
@@
As discussed extensively in Huang et al. (2022), LAEs appear to trace overdensities in a complicated manner. While certain structures appear to be well traced by LAEs, with some structures exhibiting higher levels of LAEs overdensity relative to other galaxy types, other structures appear to be largely devoid of LAEs (e.g., Apostolovski et al., 2024). In the Hyperion proto-supercluster (Cucciati et al., 2018), Huang et al. (2022) found that LAEs generally trace the overall galaxy density distribution well, with several of the most prominent matter peaks in Hyperion exhibiting strong LAE overdensities, and a significant correlation was observed between the density distribution of LAEs and spectroscopically confirmed member galaxies. However, at least one of the peaks (peak 5; Wang et al. 2016) was not well traced by LAEs, perhaps due, at least in part, to the advanced evolutionary stage of that peak. Similar variation between the density of LAEs and the overall galaxy population has also been observed in several case studies of overdensities at high redshift (e.g., Shi et al., 2019a, b)/ Additionally, it potentially originates from variations in the density of large-scale, cool H I gas reservoirs, which have the effect of diffusing or, in the presence of dust, destroying Ly photons, causing an overall suppression of LAEs. Indeed, Huang et al. (2022) also found, by cross-correlating IGM tomographic maps from the COSMOS Lyman-Alpha Mapping And Mapping Observations (CLAMATO, Lee et al. 2016, 2018) and LAEs, that while LAEs appear to trace structure moderately rich in large-scale, cool H I gas well, LAEs tend to avoid regions with the highest H I column densities (see also Momose et al. 2021). A multitude of other considerations could also modulate the overdensity of LAEs within LSSs, including the prevalence of outflows due to potential increased AGN activity (e.g., Shah et al., 2024b) or star formation activity (e.g., Lemaux et al., 2022; Staab et al., 2024), changes in the age and/or dust content of constituent member galaxies, and small number statistics due in part to LAEs being fairly rare populations at these redshifts (e.g., Cassata et al., 2015).
@@
In Figure 15 we show a comparison between ODIN-detected protoclusters and C3VO protostructure candidates in the region of COSMOS where the ODIN maps and our VMC maps overlap. We limit the C3VO protostructure candidates to those with Gaussian amplitudes 100 and a redshift range . The amplitude cut was made in order to include only those C3VO candidates with a 50% chance of being a genuine overdensity. The redshift range was chosen to order to roughly match the redshift selection function of ODIN N501 narrowband observations () with an additional () buffer in order to account for uncertainties associated with C3VO systemic redshifts and the wings of the NB501 throughput curve. In total, three ODIN protoclusters lie within the overlap volume, all three of which have at least one C3VO protostructure candidate whose barycenter falls within a projected distance of 2.5 Mpc of the ODIN protocluster barycenters. Both ODIN protoclusters in the northeast of the overlap volume appear to be potentially associated with multiple C3VO protostructure candidates at a variety of different redshifts. The ODIN protocluster at [R.A., decl.] = [150.54, 2.52] also appears to be associated with MAGAZ3NEJ1001, which is an overdensity that, perhaps not surprisingly given its presence in the ODIN map, appears to consist of primarily bluer star-forming galaxies, demographics that hold even for its most massive members (McConachie et al., 2025). All C3VO protostructure candidates that are within a proximity of 2.5 Mpc of ODIN protoclusters have moderate Gaussian amplitudes in the range of .
@@
In contrast, the inverse comparison reveals that the two most strongly detected, and presumably most massive, protostructure candidates in the C3VO map in this volume (), which are located at [R.A., decl.] = [150.25, 2.25], do not have counterpart overdensities in the ODIN maps. Additionally, several protostructure candidates with Gaussian amplitudes that are similar to those associated with the ODIN protoclusters do not have associated detections in the ODIN maps, with % of C3VO candidates lying farther than 2.5 Mpc from an ODIN protocluster. However, three out of the four of the C3VO candidates that are not associated with an ODIN protocluster lie either outside or at the edge of the redshift selection window for ODIN LAEs. The one C3VO candidate with a systemic redshift that lies well within the ODIN redshift selection window (located at [R.A., decl.] = 150.36, 2.35) has a barycenter that lies within 5 Mpc of an ODIN-selected protocluster in projection. Furthermore, the two strongest C3VO detections in this volume were detected as potential overdensities in Ramakrishnan et al. (2023), but did not exhibit a strong enough overdensity over a sufficiently large area to pass their fiducial cuts for classification as a genuine protocluster. As both of these C3VO candidates lie at the extrema of the ODIN redshift selection window, it is likely that the ODIN observations are only able to recover a small fraction of member galaxies in these systems.
@@
Structure | R.A.lit | Decl.lit | R.A.C3VO | Decl.C3VO | Amp.aaAmplitude of the Gaussian fit to the linked candidate. | dtrans | IDC3VOb | ref. | ||
@@
C.PS4 | 150.4500 | 2.3710 | 2.035 | 150.4395 | 2.3726 | 2.024 | 265 | 0.33 | C_SP38 | 1,2 |
COSTCO2 | 149.8710 | 2.2290 | 2.047 | 149.9323 | 2.1311 | 2.089 | 162 | 3.57∗ | C_SP93 | 3 |
ZFIRE | 150.0940 | 2.2510 | 2.095 | 150.0459 | 2.2415 | 2.112 | 322 | 1.51+ | C_SP28 | 4 |
G237 | 150.5070 | 2.3120 | 2.160 | 150.4889 | 2.3617 | 2.146 | 237 | 1.62 | C_SP54 | 5,6 |
COSTCO3 | 150.1290 | 2.2750 | 2.160 | 150.0957 | 2.2449 | 2.208 | 284 | 1.38 | C_SP33 | 3 |
C.PS6 | 150.1900 | 2.1330 | 2.174 | 150.1786 | 2.1887 | 2.153 | 166 | 1.74 | C_SP87 | 1,2 |
C.PS8 | 149.8640 | 2.0900 | 2.216 | 149.8603 | 2.0950 | 2.215 | 615 | 0.19 | C_SP7 | 1,2 |
BD2.2 | 150.1970 | 2.0030 | 2.232 | 150.1848 | 2.0176 | 2.226 | 929 | 0.58+ | C_SP6 | 7 |
C.PS9 | 149.9200 | 2.2210 | 2.277 | 149.8851 | 2.1652 | 2.284 | 455 | 2.00 | C_SP14 | 1,2,8 |
COSTCO5 | 149.9380 | 2.0910 | 2.283 | 150.0112 | 2.1181 | 2.262 | 98 | 2.38 | C_SP175 | 3 |
COSTCO1 | 150.1100 | 2.1610 | 2.298 | 150.1021 | 2.2218 | 2.301 | 380 | 1.86 | C_SP34 | 3 |
COSTCO4 | 149.7060 | 2.0240 | 2.391 | 149.7715 | 2.1389 | 2.416 | 305 | 3.99∗,+ | C_SP29 | 3 |
Theia | 150.0937 | 2.4049 | 2.468 | 150.0951 | 2.4086 | 2.468 | 988 | 0.12+ | C_SP5 | 9, 10, 11 |
Eos | 149.9765 | 2.1124 | 2.426 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.31 | C_SP2 | 9 |
Helios | 149.9996 | 2.2537 | 2.444 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.25 | C_SP2 | 8, 9, 12, 13 |
Selene | 150.2556 | 2.3423 | 2.469 | 150.1992 | 2.3781 | 2.462 | 2013 | 2.00+ | C_SP1 | 9, 11 |
Hyp5 | 150.2293 | 2.3381 | 2.507 | 150.1992 | 2.3781 | 2.462 | 2013 | 1.50+ | C_SP1 | 9, 14 |
Hyp6 | 150.3316 | 2.2427 | 2.492 | 150.3403 | 2.2236 | 2.496 | 370 | 0.63+ | C_SP22 | 9 |
Hyp7 | 149.9581 | 2.2187 | 2.423 | 150.0125 | 2.1124 | 2.432 | 1640 | 2.01 | C_SP2 | 9 |
LATIS1-D2-0 | 150.1030 | 2.1750 | 2.562 | 150.1327 | 2.3380 | 2.558 | 417 | 4.93∗,+ | C_SP18 | 15 |